OpenStudy (anonymous):

Given the linear equation y = 4x + 2, find the y-coordinates of the points (-2, ), (1, ), and (-3, ). Please help? I am not understanding solving equations pertaiing to graphs.

myininaya (myininaya):

just plug in what is y if x=-2 y=4(-2)+2=-8+2=-6 (-2,-6)

OpenStudy (anonymous):

Plug in the "x-values" given, and solve for the y value.

myininaya (myininaya):

you can do the rest

OpenStudy (anonymous):

what do I need to do with -2 1 & -3

OpenStudy (anonymous):

Plug those values in for x in the y=4x+2 equation. myininaya gave an example above.

OpenStudy (anonymous):

so use three different equations x =-2, x=1, x=-3?

myininaya (myininaya):

use the same equation y=4x+2 (-2, ___) -2 is your input find out what your output y is put pluggin in -2 where that x is y=4(-2)+2=-8+2=-6 (-2,-6)

OpenStudy (anonymous):

i got that part. I am asking to also plug 1 and -3 in for the place of x?

myininaya (myininaya):

ok just do as i did above

myininaya (myininaya):

just put that 1 where x is and evaluate y